# Phase space analysis of CCDM cosmologies

### v1

@@ III.5.2 Two fluid system
$F_{3}(1,1)$ Always Always Saddle $1$ Always yes
$F_{4}\left(1,\frac{3-\beta}{3-\beta+\alpha}\right)$ $\beta\leq 3$ Stable if $w>-1$ $1$ Always yes
$F_{5}\left(0,z_{c}\right)$ $w=-1$ Stable if $\beta-3+(3-\beta+\alpha)z_{c}>0$ 00 Always yes
@@
$F_{7}\left(x_{c},0\right)$ $03w+\beta=$ 0 Stable if $w<-1$ $x_{c}$ Yes if $\beta>1$
$\Gamma_{0}+\Gamma_{3}H^{-1}$ $G_{0}(0,0)$ Always Stable if $w<-1$; 00 Yes if $w<-\frac{1}{3}$
$G_{1}(1,0)$ Always Saddle if $w>0$, Unstable if $w<0$; $1$ No
$G_{2}(0,1)$ Always Unstable/Saddle 00 Undetermined
$G_{3}(1,1)$ Always Saddle $1$ Always yes
$G_{4}\left(1,\frac{6}{6+\alpha+\sqrt{\alpha^{2}+12\mu}}\right)$ Always Always Stable $1$ Always yes
@@
$G_{6}\left(x_{c},\frac{6}{6+\alpha+\sqrt{\alpha^{2}+12\mu}}\right)$ $w=-1$ Always Stable $x_{c}$ Always yes
$G_{7}\left(0,z_{c}\right)$ $w=-1$ Stable if $z_{c}>(<)\frac{6}{6+\alpha+\sqrt{\alpha^{2}+12\mu}}$, $03+\alpha-\mu>(<)$ 0; 00 Always yes
Unstable if $z_{c}<(>)\frac{6}{6+\alpha+\sqrt{\alpha^{2}+12\mu}}$, $03+\alpha-\mu>(<)$ 0
$\Gamma_{0}+\Gamma_{2}H^{2}$ $I_{0}(0,0)$ Always Unstable/Saddle 00 Undetermined
$I_{1}(1,0)$ Always Always Stable $1$ Always yes
$I_{2}(0,1)$ Always Unstable/Saddle 00 Undetermined
$I_{3}(1,1)$ Always Saddle $1$ Always yes
$I_{4\pm}\left(1,\frac{3+2\gamma\pm\sqrt{9-4\alpha\gamma}}{2(3+\alpha+\gamma)}\right)$ $4\alpha\gamma\leq 9$ $I_{4+}:$ Stable if $w>-1$, otherwise Saddle $1$ Always yes
@@
$I_{5\pm}\left(x_{c},\frac{3+2\gamma\pm\sqrt{9-4\alpha\gamma}}{2(3+\alpha+\gamma)}\right)$ $14\alpha\gamma\leq 9,~{}w=-$ 1 $I_{5+}:$ Stable; $I_{5-}:$ Unstable $x_{c}$ Always yes
$I_{6}\left(0,z_{c}\right)$ $w=-1$ Stable if $\frac{3+2\gamma-\sqrt{9-4\alpha\gamma}}{2(3+\alpha+\gamma)}<z_{c}<\frac{3+2\gamma+\sqrt{9-4\alpha\gamma}}{2(3+\alpha+\gamma)}$ 00 Always yes
@@
Two-fluid system
Creation rate Critical point Existence Stability $\Omega_{dm}$ Acceleration
$\Gamma_{1}H+\Gamma_{2}H^{2}$ $J_{0}(0,0)$ Always Unstable/Saddle 00 Undetermined
$J_{1}(1,0)$ Always Always Stable $1$ Always yes
$J_{2}(0,1)$ Always Stable if $w\leq-\frac{\beta}{3}$, $w>-1$ 00 Yes if $w<-\frac{1}{3}$
$J_{3}(1,1)$ Always Stable if $w>-\frac{\beta}{3}$, $\beta<3$ $1$ Yes if $\beta>1$
$J_{4}\left(1,\frac{\gamma}{\gamma+3-\beta}\right)$ $\beta\leq 3$ Saddle if $w>-1$; Unstable if $w<-1$ $1$ Always yes
@@
$J_{6}\left(0,z_{c}\right)$ $w=-1$ Stable if $\gamma-(\gamma+3-\beta)z_{c}<0$ 00 Always yes
@@
$\Gamma_{1}H+\Gamma_{3}H^{-1}$ $L_{0}(0,0)$ Always Stable if $w\leq-\frac{\beta}{3},~{}w<-1$ 00 Yes if $w<-\frac{1}{3}$
$L_{1}(1,0)$ Always Stable if $w>-\frac{\beta}{3},~{}\beta\geq 3$ $1$ Yes if $\beta>1$
$L_{2}(0,1)$ Always Unstable/Saddle 00 Undetermined
$L_{3}(1,1)$ Always Saddle $1$ Always yes
$L_{4}\left(1,\frac{3-\beta-\sqrt{(3-\beta)\mu}}{3-\beta-\mu}\right)$ $\beta\leq 3$ Stable if $w>-1,~{}\beta+\mu\neq 3$ $1$ Always yes
@@
$L_{6}\left(0,z_{c}\right)$ $w=-1$ Stable if $z_{c}<\frac{3-\beta-\sqrt{(3-\beta)\mu}}{3-\beta-\mu}$; Unstable if $z_{c}>\frac{3-\beta-\sqrt{(3-\beta)\mu}}{3-\beta-\mu}$ 00 Always yes
@@
$\Gamma_{0}+\Gamma_{1}H+\Gamma_{2}H^{2}$ $M_{0}(0,0)$ Always Unstable/Saddle 00 Undetermined
$M_{1}(1,0)$ Always Always Stable $1$ Always yes
$M_{2}(0,1)$ Always Unstable/Saddle 00 Undetermined
$M_{3}(1,1)$ Always Saddle $1$ Always yes
$M_{4\pm}\left(1,\frac{3-\beta+2\gamma\pm\sqrt{(3-\beta)^{2}-4\alpha\gamma}}{2(3+\alpha-\beta+\gamma)}\right)$ $(3-\beta)^{2}-4\alpha\gamma\geq 0$, $M_{4+}:$ Stable if $w>-1$, Saddle if $w<-1$ $1$ Always yes
$3+2\alpha>\beta,~{}3+2\gamma>\beta$; $M_{4-}:$ Unstable if $w<-1$, Saddle if $w>-1$
@@
$M_{6}\left(0,z_{c}\right)$ $w=-1$ Stable if $\frac{3-\beta+2\gamma-\sqrt{(3-\beta)^{2}-4\alpha\gamma}}{2(3+\alpha-\beta+\gamma)}<z_{c}<\frac{3-\beta+2\gamma+\sqrt{(3-\beta)^{2}-4\alpha\gamma}}{2(3+\alpha-\beta+\gamma)}$, 00 Always yes

### v2

@@ VI.6 Model: $\Gamma=\Gamma_{0}+\Gamma_{1}H+\Gamma_{2}H^{2}$
$F_{3}(1,1)$ Always Always Saddle 11 Always yes
$F_{4}\left(1,\frac{3-\beta}{3-\beta+\alpha}\right)$ $\beta\leq 3$ Stable if w>1w>-1 11 Always yes
$F_{5}\left(0,z_{c}\right)$ w=1w=-1 Stable if $\beta-3+(3-\beta+\alpha)z_{c}>0$ 0 Always yes
@@
$F_{7}\left(x_{c},0\right)$ $3w+\beta=0$ Stable if w<1w<-1 $x_{c}$ Yes if $\beta>1$
$\Gamma_{0}+\Gamma_{3}H^{-1}$ $G_{0}(0,0)$ Always Stable if w<1w<-1; 0 Yes if $w<-\frac{1}{3}$
$G_{1}(1,0)$ Always Saddle if $w>0$, Unstable if $w<0$; 11 No
$G_{2}(0,1)$ Always Unstable/Saddle 0 Undetermined
$G_{3}(1,1)$ Always Saddle 11 Always yes
$G_{4}\left(1,\frac{6}{6+\alpha+\sqrt{\alpha^{2}+12\mu}}\right)$ Always Always Stable 11 Always yes
@@
$G_{6}\left(x_{c},\frac{6}{6+\alpha+\sqrt{\alpha^{2}+12\mu}}\right)$ w=1w=-1 Always Stable $x_{c}$ Always yes
$G_{7}\left(0,z_{c}\right)$ w=1w=-1 Stable if $z_{c}>(<)\frac{6}{6+\alpha+\sqrt{\alpha^{2}+12\mu}}$, $3+\alpha-\mu>(<)0$; 0 Always yes
Unstable if $z_{c}<(>)\frac{6}{6+\alpha+\sqrt{\alpha^{2}+12\mu}}$, $3+\alpha-\mu>(<)0$
$\Gamma_{0}+\Gamma_{2}H^{2}$ $I_{0}(0,0)$ Always Unstable/Saddle 0 Undetermined
$I_{1}(1,0)$ Always Always Stable 11 Always yes
$I_{2}(0,1)$ Always Unstable/Saddle 0 Undetermined
$I_{3}(1,1)$ Always Saddle 11 Always yes
$I_{4\pm}\left(1,\frac{3+2\gamma\pm\sqrt{9-4\alpha\gamma}}{2(3+\alpha+\gamma)}\right)$ $4\alpha\gamma\leq 9$ $I_{4+}:$ Stable if w>1w>-1, otherwise Saddle 11 Always yes
@@
$I_{5\pm}\left(x_{c},\frac{3+2\gamma\pm\sqrt{9-4\alpha\gamma}}{2(3+\alpha+\gamma)}\right)$ $4\alpha\gamma\leq 9,~w=-1$ $I_{5+}:$ Stable; $I_{5-}:$ Unstable $x_{c}$ Always yes
$I_{6}\left(0,z_{c}\right)$ w=1w=-1 Stable if $\frac{3+2\gamma-\sqrt{9-4\alpha\gamma}}{2(3+\alpha+\gamma)}<z_{c}<\frac{3+2\gamma+\sqrt{9-4\alpha\gamma}}{2(3+\alpha+\gamma)}$ 0 Always yes
@@
Creation rate Critical point Existence Stability $\mathbf{\Omega_{dm}^{\rm eff}}$ Acceleration
$\Gamma_{1}H+\Gamma_{2}H^{2}$ $J_{0}(0,0)$ Always Unstable/Saddle 0 Undetermined
$J_{1}(1,0)$ Always Always Stable 11 Always yes
$J_{2}(0,1)$ Always Stable if $w\leq-\frac{\beta}{3}$, w>1w>-1 0 Yes if $w<-\frac{1}{3}$
$J_{3}(1,1)$ Always Stable if $w>-\frac{\beta}{3}$, $\beta<3$ 11 Yes if $\beta>1$
$J_{4}\left(1,\frac{\gamma}{\gamma+3-\beta}\right)$ $\beta\leq 3$ Saddle if w>1w>-1; Unstable if w<1w<-1 11 Always yes
@@
$J_{6}\left(0,z_{c}\right)$ w=1w=-1 Stable if $\gamma-(\gamma+3-\beta)z_{c}<0$ 0 Always yes
@@
$\Gamma_{1}H+\Gamma_{3}H^{-1}$ $L_{0}(0,0)$ Always Stable if $w\leq-\frac{\beta}{3},~w<-1$ 0 Yes if $w<-\frac{1}{3}$
$L_{1}(1,0)$ Always Stable if $w>-\frac{\beta}{3},~\beta\geq 3$ 11 Yes if $\beta>1$
$L_{2}(0,1)$ Always Unstable/Saddle 0 Undetermined
$L_{3}(1,1)$ Always Saddle 11 Always yes
$L_{4}\left(1,\frac{3-\beta-\sqrt{(3-\beta)\mu}}{3-\beta-\mu}\right)$ $\beta\leq 3$ Stable if $w>-1,~\beta+\mu\neq 3$ 11 Always yes
@@
$L_{6}\left(0,z_{c}\right)$ w=1w=-1 Stable if $z_{c}<\frac{3-\beta-\sqrt{(3-\beta)\mu}}{3-\beta-\mu}$; Unstable if $z_{c}>\frac{3-\beta-\sqrt{(3-\beta)\mu}}{3-\beta-\mu}$ 0 Always yes
@@
$\Gamma_{0}+\Gamma_{1}H+\Gamma_{2}H^{2}$ $M_{0}(0,0)$ Always Unstable/Saddle 0 Undetermined
$M_{1}(1,0)$ Always Always Stable 11 Always yes
$M_{2}(0,1)$ Always Unstable/Saddle 0 Undetermined
$M_{3}(1,1)$ Always Saddle 11 Always yes
$M_{4\pm}\left(1,\frac{3-\beta+2\gamma\pm\sqrt{(3-\beta)^{2}-4\alpha\gamma}}{2(3+\alpha-\beta+\gamma)}\right)$ $(3-\beta)^{2}-4\alpha\gamma\geq 0$, $M_{4+}:$ Stable if w>1w>-1, Saddle if w<1w<-1 11 Always yes
$3+2\alpha>\beta,~3+2\gamma>\beta$; $M_{4-}:$ Unstable if w<1w<-1, Saddle if w>1w>-1
@@
$M_{6}\left(0,z_{c}\right)$ w=1w=-1 Stable if $\frac{3-\beta+2\gamma-\sqrt{(3-\beta)^{2}-4\alpha\gamma}}{2(3+\alpha-\beta+\gamma)}<z_{c}<\frac{3-\beta+2\gamma+\sqrt{(3-\beta)^{2}-4\alpha\gamma}}{2(3+\alpha-\beta+\gamma)}$, 0 Always yes